# Elimination Without Eliminating: Computing Complements of Real Hypersurfaces Using Pseudo-Witness Sets

### v1

@@ 3.2 Evaluating the gradient and the Hessian
We now turn to the problem of evaluating $\nabla r$ and ${\rm hess}(r)$ of the unknown defining polynomial $h$ of $\mathcal{H}\subset{\mathbb{C}}^{k}$, using the data of a pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ where $\mathcal{L}\subset{\mathbb{C}}^{k}$ is a line. We underline that the pseudo-witness set need not be reduced here. The pseudo-witness set in this section is only used to give access to the points in $\mathcal{H}\cap\mathcal{L}$.
It is described in [22] how to utilize a pseudo-witness set to evaluate and differentiate $h$ along the line $\mathcal{L}$. This implies that $\nabla h(p)$ can be computed from $k$ linearly independent lines $\mathcal{L}$ through $p$. Likewise, ${\rm hess}(h)(p)$ can be computed through $k^{2}$ lines through $p$. In this section, we prove that, in fact, a single general line $\mathcal{L}$ through $p$ suffices to determine both $\nabla h(p)$ and ${\rm hess}(h)(p)$. This drastically improves the complexity of the computation.
@@
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\Bigl(\log|h(p+tb)|\Bigr)\Big|_{t=0}&=\frac{\mathrm{d}}{\mathrm{d}t}\Biggl(\sum\limits_{j=1}^{\deg h}\log|t-t_{j}|\Biggr)\Biggr|_{t=0}=\sum\limits_{j=1}^{\deg h}\frac{\mathrm{d}}{\mathrm{d}t}\Big(|t-t_{j}|\Big)\Bigr|_{t=0}=-\sum_{j=1}^{\deg h}\frac{1}{t_{j}}.\end{split}$
@@
$\nabla_{p}\log|h(p)|=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)\qquad\text{and}\qquad{\rm hess}(\log|h(p)|)=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr).$
@@
Output: The gradient $\nabla\log|h|$ and the Hessian ${\rm hess}(\log|h|)$ evaluated at $p$.
@@
Solving this system yields $\nabla_{b}(\nabla_{p}s)\in{\mathbb{C}}^{2\times 2}$. Summing over all $s_{j}$ we obtain, according to the formula in Theorem 3.8, ${\rm hess}(\log|h|)(p)=-\nabla_{b}(\nabla_{p}s_{1})-\nabla_{b}(\nabla_{p}s_{2})$.
@@
$\nabla_{p}\log r(p)$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)-\frac{2(p-c)}{q(p)},\quad\text{and}$
${\rm hess}(\log r(p))$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr)+\frac{4}{q(p)^{2}}(p-c)(p-c)^{T}-\frac{2}{q(p)}I_{n}.$
@@ 3.3 Summary of the algorithm for unknown defining polynomial
We summarize our strategy for computing complements of hypersurfaces defined by pseudo-witness sets in Algorithm 4. An implementation of the algorithm in Julia, built on top of the numerical algebraic geometry package HomtopyContinuation.jl [10] will be provided in a forthcoming software paper. In the next section, we showcase the result of three early proof-of-concept calculations from this implementation.
@@
(3) Compute routing points. Compute all solutions to $\nabla\log r(x)=0$ by a two step homotopy: First sample a random $q\in\mathbb{C}^{k}$ and compute the all complex solutions of $\nabla\log r(x)-q=0$ by varying $q$ in monodromy loops. Then, track those solutions along the homotopy $H(x,t)=\nabla\log r(x)-tq$ for $t$ from $1$ to 0. Filter out the real solutions. These are the routing points.
(4) Compute regions. Run steps (2)–(5) in Algorithm 1.
@@
For Algorithm 4, we do not need to assume that pseudo-witness sets are reduced as long as we have a method to compute the points in $\mathcal{H}\cap\mathcal{L}$. However, when using numerical homotopy continuation for computing $\mathcal{H}\cap\mathcal{L}$ when evaluating $\nabla r$ and ${\rm hess}(r)$, we must assume reducedness. Suppose that $X_{1}\ldots,X_{\ell}$ are the irreducible components of $X$ such that, after reordering, the first $j$ are reduced, i.e., $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},X_{i}\cap(\mathcal{L}\times\mathcal{L}^{\prime}))$ for $1\leq i\leq j\leq\ell$ are reduced pseudo-witness sets. We can track these pseudo-witness sets towards other linear spaces using homotopy continuation, since only for reduced pseudo-witness sets the second condition in (2.2) is satisfied. Consequently, we can run Algorithm 4 only on the hypersurface $\hat{\mathcal{H}}=\pi(X_{1}\cup\cdots\cup X_{j})\subset\mathcal{H}$, but not on $\mathcal{H}$. We will see an example of this in Section 4.3. This is an inherent limitation of our algorithm when using numerical homotopy continuation, which is the only method fast enough to compute the numerous sequential intersections with linear spaces that are required for each evaluation of $\nabla\log|h|$ and ${\rm hess}(\log|h|)$ via Algorithm 3.
The computational bottleneck in Algorithm 4 is solving the system $\nabla\log r=0$ with monodromy, as each monodromy loop requires many evaluations of $\nabla\log|h|$ and ${\rm hess}(\log|h|)$. To speed up this step, one can extend the set of start solutions as follows. We sample some number of points in ${\mathbb{R}}^{k}$, and use (1) gradient flow to trace them to routing points (of index 0), which are solutions of $\nabla\log r=0$ and (2) Newton method on $\nabla\log r$, which, if it converges, return solutions to $\nabla\log r=0$ We then use a parameter homotopy to trace them to additional start solutions of $\nabla\log r-q_{0}=0$.
@@ 4 Case studies
In this section, we illustrate our method on three polynomial systems arising in applications: the Kuramoto model (Section 4.1), the 3RPR mechanism from kinematics (Section 4.2), and a model for the Allee effect in population dynamics (Section 4.3). Each example consists of a parameterized polynomial system
@@
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection to the parameter space, and overline denotes the Zariski closure. The real complement ${\mathbb{R}}^{k}\smallsetminus\Delta$ has finitely many regions. In each of these regions, the system has a constant, finite number of roots. In what follows, we compute the regions using the method outlined in Algorithm 4.
@@ 4.1 Kuramoto model
where $K$ is the coupling strength, $N$ is the number of oscillators, $\Omega=(\omega_{1},\dots,\omega_{N})$ is the vector of frequencies, and $a_{i,j}\in\{0,1\}$ is the $(i,j)$-th element of the adjacency matrix of the coupling graph.
Fixing $\theta_{N}=0$ and using the identity $\sin(x-y)=\sin(x)\cos(y)-\sin(y)\cos(x)$ we substitute $s_{i}=\sin(\theta_{i})$ and $c_{i}=\cos(\theta_{i})$ to rewrite the remaining N1N-1 equations into polynomial form. Coupling these equations with the Pythagorean identity $s_{i}^{2}+c_{i}^{2}=1$, we obtain a system of 2(N1)2(N-1) polynomial equations in 2(N1)2(N-1) variables:
$\begin{split}0&=\omega_{i}+\frac{K}{N}\sum\limits_{j=1}^{N}a_{i,j}(s_{i}c_{j}-s_{j}c_{i})\\ 0&=s_{i}^{2}+c_{i}^{2}-1.\end{split}$ (4.3)
We study the Kuramoto model with three oscillators, where the graph is the triangle; i.e., $a_{i,j}=1$ whenever $i\neq j$ and 0 otherwise. In the formulation of (4.3), this gives rise to the polynomial system
$G(\omega;z)=\begin{pmatrix}(s_{1}c_{2}-c_{1}s_{2})+(s_{1}c_{3}-c_{1}s_{3})-3\omega_{1}\\[1.99997pt] (s_{2}c_{1}-c_{2}s_{1})+(s_{2}c_{3}-c_{2}s_{3})-3\omega_{2}\\[1.99997pt] s_{1}^{2}+c_{1}^{2}-1\\[1.99997pt] s_{2}^{2}+c_{2}^{2}-1\end{pmatrix}$ (4.4)
@@
Figure 4: Routing points and connecting paths for the discriminant of the two-parameter version of the Kuramoto model with three oscillators (4.4). The left picture shows a zoomed out version of the right picture. The gradient flow is visible in blue. The arrows point towards index-zero routing points. We see that 77 out of 99 regions are inside the star-shaped figure. The last two regions are the inside of the ellipse and the unbounded region outside the ellipse. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
@@
To model its motion, two coordinate frames are used: a fixed frame attached to the ground, where the anchor points lie at fixed positions, and a moving frame attached to the triangular platform, where the platform’s corner points remain fixed relative to each other. Each leg connects one anchor point to one platform corner. In Figure 5, the leg lengths are $\ell_{1},\ell_{2}$, and $\ell_{3}$, and we work with their squares $c_{i}=\ell_{i}^{2}$ for convenience. In the fixed frame, the anchor points are located at $(0,0)$, $(A_{2},0)$, and $(A_{3},B_{3})$ while in the moving frame, the platform corners are located at the points $P_{1}=(0,0)$, $P_{2}=(a_{2},0)$, and $P_{3}=(a_{3},b_{3})$.
@@
A fundamental problem is to determine all the ways a real motion of the mechanism can start and end in the same “home” configuration for a fixed set of leg lengths. The configuration variables are $(p,\phi)=(p_{1},p_{2},\phi_{1},\phi_{2})$, where $p=(p_{1},p_{2})$ represents the translation of the platform and $\phi=(\phi_{1},\phi_{2})$ is a unit circle representation of its rotation. This yields a polynomial system $G(a,A,b_{3},B,c;p,\phi)$ with variables $(p,\phi)$, parameters
@@
Figure 6: Routing points and connecting paths for the two-parameter version of the 3RPR system. The left picture shows a zoomed out version of the right picture. The gradient flow is visible in blue. The arrows point towards index-zero routing points. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
@@
Finally, we consider a three-parameter version of the problem where we instead view $(c_{1},c_{2},A_{2})$ as free parameters, and fix other parameters as in (4.5). In this case, we have not been able to find a symbolic expression for the discriminant surface, but its degree can readily be determined to 24 through a pseudo witness set. This example is at the limit of what our current implementation can handle, as the monodromy step did not terminate after 14 days [^1] Computations were run on a 4-socket Intel Xeon Gold 6128 system (24 cores total, 3.4 GHz) with 3 TB RAM.. At this point, 457 complex critical points for the routing function centered at $(4.72,4.33,1.70)$ had been found, of which 90 points were real. While likely not being a complete set of routing points, these points still provide a valuable sample of the parameter space. In particular, we can use them to verify that there are open regions in parameter space with the real root counts 2, 4 and 6, respectively.
@@ 4.3 Allee effect
To this end, we extend the discriminant $\Delta$ from (4.1) to $\Delta\cup V(ab(b-\frac{1}{2}))$. As explained in Remark 3.1, on the level of routing functions, this means adding $\log(|a|)$, $\log(|b|)$ and $\log(|b-\frac{1}{2}|)$ to $\log(r)$. It is clear from inspection of $G$ that $z=0$ is always a solution for biologically relevant parameters, and that all other nonnegative solutions have strictly positive coordinates. Hence, the number of nonnegative solutions is constant in each region of $((0,\infty)\times(0,\frac{1}{2}))\smallsetminus(\Delta\cup V(ab(b-\frac{1}{2}))$.
@@
An interesting property of this example is that $\Delta$ is not irreducible, and that some of its irreducible components have a non-reduced pseduo-witness set with respect to the equations in  (4.1). Hence, if these equations are used, our implementation effectively works with the subvariety $\hat{\Delta}\subset\Delta$ corresponding to the reduced part (see Remark 3.11). We can see the effect of this in the left part of Figure 7, which shows the resulting routing function, the biologically relevant routing points and the number of nonnegative solutions of $G(a,b,z)=0$ inside the green critical points. As we can see, gradient flow connects a parameter pair $(a,b)$, such that $G(a,b,z)=0$ has 15 positive solutions with another pair, for which there are 9 positive solutions.
Figure 7: Contours and critical points of a routing functions for the modified discriminant $\Delta\cup V(ab(b-\frac{1}{2}))$ of the Allee system from Section 4.3. The left figure corresponds to using the formulation (4.1), which gives the subvariety $\hat{\Delta}$. The right figure uses the radical and gives the full discriminant $\Delta$. The routing points are in green and are labeled by the number of nonnegative solutions of $G(a,b,z)=0$. In both figures, there is a routing point of the unbounded region which is outside the picture, with nonnegative root count 3. Contours (up to a constant), critical points and gradient flow were computed without access to the discriminant polynomial.
@@
$I=\left\langle G(a,b,z),\ \tfrac{\partial G}{\partial z}(a,b,z)\right\rangle$
can readily be computed symbolically in the computer algebra system OSCAR [18]. By using a generating set of $\sqrt{I}$, together with the standard “squaring up” technique [4, Section 2.5] for dealing with overdetermined systems, we obtain a reduced pseudo-witness set for $\Delta$. The resulting routing function and the biologically relevant routing points are shown in the right part of Figure 7, which agrees with the regions displayed in [44, Figure 3].

### v2

@@ 3.2 Evaluating the gradient and the Hessian
We now turn to the problem of evaluating $\nabla r$ and $\operatorname*{Hess}(r)$ of the unknown defining polynomial $h$ of $\mathcal{H}\subset{\mathbb{C}}^{k}$, using the data of a pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ where $\mathcal{L}\subset{\mathbb{C}}^{k}$ is a line. We underline that the pseudo-witness set need not be reduced here. The pseudo-witness set in this section is only used to give access to the points in $\mathcal{H}\cap\mathcal{L}$.
It is described in [22] how to utilize a pseudo-witness set to evaluate and differentiate $h$ along the line $\mathcal{L}$. This implies that $\nabla h(p)$ can be computed from $k$ linearly independent lines $\mathcal{L}$ through $p$. Likewise, $\operatorname*{Hess}(h)(p)$ can be computed through $k^{2}$ lines through $p$. In this section, we prove that, in fact, a single general line $\mathcal{L}$ through $p$ suffices to determine both $\nabla h(p)$ and $\operatorname*{Hess}(h)(p)$. This drastically improves the complexity of the computation.
@@
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\Bigl(\log|h(p+tb)|\Bigr)\Big|_{t=0}&=\frac{\mathrm{d}}{\mathrm{d}t}\Biggl(\sum\limits_{j=1}^{\deg h}\log|t-t_{j}|\Biggr)\Biggr|_{t=0}=\sum\limits_{j=1}^{\deg h}\frac{\mathrm{d}}{\mathrm{d}t}\log\Big(|t-t_{j}|\Big)\Bigr|_{t=0}=-\sum_{j=1}^{\deg h}\frac{1}{t_{j}}.\end{split}$
@@
$\nabla_{p}\log|h(p)|=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)\qquad\text{and}\qquad\operatorname*{Hess}(\log|h(p)|)=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr).$
@@
Output: The gradient $\nabla\log|h|$ and the Hessian $\operatorname*{Hess}(\log|h|)$ evaluated at $p$.
@@
Solving this system yields $\nabla_{b}(\nabla_{p}s)\in{\mathbb{C}}^{2\times 2}$. Summing over all $s_{j}$ we obtain, according to the formula in Theorem 3.8, $\operatorname*{Hess}(\log|h|)(p)=-\nabla_{b}(\nabla_{p}s_{1})-\nabla_{b}(\nabla_{p}s_{2})$.
@@
$\nabla_{p}\log r(p)$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)-\frac{2e(p-c)}{q(p)},\quad\text{and}$
$\operatorname*{Hess}(\log r(p))$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr)+\frac{4e}{q(p)^{2}}(p-c)(p-c)^{T}-\frac{2e}{q(p)}I_{k}.$
@@ 3.3 Summary of the algorithm for unknown defining polynomial
We summarize our strategy for computing complements of hypersurfaces defined by pseudo-witness sets in Algorithm 4. An implementation of the algorithm in Julia, built on top of the numerical algebraic geometry package HomotopyContinuation.jl [10] will be provided in a forthcoming software paper. In the next section, we showcase the result of three early proof-of-concept calculations from this implementation.
@@
(3) Compute routing points. Compute all solutions to $\nabla\log r(p)=0$ by a two-step homotopy: First sample a random $q\in\mathbb{C}^{k}$ and compute all complex solutions of $\nabla\log r(p)-q=0$ by varying $q$ in monodromy loops. Then, track those solutions along the homotopy $H(p,t)=\nabla\log r(p)-tq$ for $t$ from $1$ to 0. Filter out the real solutions.
(4) Compute regions. Run Algorithm 1.
@@
For Algorithm 4, we do not need to assume that pseudo-witness sets are reduced as long as we have a method to compute the points in $\mathcal{H}\cap\mathcal{L}$. However, when using numerical homotopy continuation for computing $\mathcal{H}\cap\mathcal{L}$ when evaluating $\nabla r$ and $\operatorname*{Hess}(r)$, we must assume reducedness. Suppose that $X_{1},\ldots,X_{\ell}$ are the irreducible components of $X$ such that, after reordering, the first $j$ are reduced, i.e., $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},X_{i}\cap(\mathcal{L}\times\mathcal{L}^{\prime}))$ for $1\leq i\leq j\leq\ell$ are reduced pseudo-witness sets. We can track these pseudo-witness sets towards other linear spaces using homotopy continuation, since only for reduced pseudo-witness sets the second condition in (2.2) is satisfied. Consequently, we can run Algorithm 4 only on the hypersurface $\hat{\mathcal{H}}=\pi(X_{1}\cup\cdots\cup X_{j})\subset\mathcal{H}$, but not on $\mathcal{H}$. We will see an example of this in Section 4.3. This is an inherent limitation of our algorithm when using numerical homotopy continuation, which is the only method fast enough to compute the numerous sequential intersections with linear spaces that are required for each evaluation of $\nabla\log|h|$ and $\operatorname*{Hess}(\log|h|)$ via Algorithm 3.
The computational bottleneck in Algorithm 4 is solving the system $\nabla\log r=0$ with monodromy, as each monodromy loop requires many evaluations of $\nabla\log|h|$ and $\operatorname*{Hess}(\log|h|)$. To speed up this step, one can extend the set of start solutions by the following two strategies: (1) sample points in ${\mathbb{R}}^{k}$ and apply gradient flow with respect to $\nabla\log r$ to find routing points of index 0, and then trace these points to solutions of $\nabla\log r-q_{0}=0$ via a parameter homotopy from $\nabla\log r=0$; (2) sample points in ${\mathbb{C}}^{k}$ and apply Newton’s method with respect to the system $\nabla\log r-q_{0}=0$.
@@ 4 Case studies
We end the paper by using our method to analyze the possible root counts of three parametric polynomial systems arising in applications: the Kuramoto model (Section 4.1), the 3RPR mechanism from kinematics (Section 4.2), and a model for the Allee effect in population ecology (Section 4.3). Each example consists of a parametric polynomial system
@@
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection to the parameter space, and overline denotes the Zariski closure. The real complement ${\mathbb{R}}^{k}\smallsetminus\Delta$ has finitely many regions. In each region, the system has a constant, finite number of roots. In what follows, we compute the regions using the method outlined in Algorithm 4.
@@ 4.1 Kuramoto model
where $\theta=(\theta_{1},\ldots,\theta_{N})$ are the phases of the oscillators, $K$ is the coupling strength, $\omega=(\omega_{1},\dots,\omega_{N})$ are the natural frequencies, and $A=(a_{i,j})$ is the adjacency matrix of the coupling graph.
Fixing $\theta_{N}=0$ and using the identity $\sin(x-y)=\sin(x)\cos(y)-\sin(y)\cos(x)$ we substitute $s_{i}=\sin(\theta_{i})$ and $c_{i}=\cos(\theta_{i})$ to rewrite the remaining N1N-1 equations into polynomial form. Coupling these equations with the Pythagorean identity $s_{i}^{2}+c_{i}^{2}=1$, we obtain a system of 2(N1)2(N-1) polynomial equations in 2(N1)2(N-1) variables, where $c_{N}=1$ and $s_{N}=0$:
$\begin{split}0=\omega_{i}+\frac{K}{N}\sum\limits_{j=1}^{N}a_{i,j}(s_{i}c_{j}-s_{j}c_{i}),\qquad 0=s_{i}^{2}+c_{i}^{2}-1,\qquad\text{for $i=1,\ldots,{N-1}$}\end{split}$ (4.3)
We study the Kuramoto model with three oscillators, where the graph is the triangle; i.e., $a_{i,j}=1$ if $i\neq j$ and 0 otherwise. In the formulation of (4.3), this gives rise to the system
$G(\omega;z)=\begin{pmatrix}(s_{1}c_{2}-c_{1}s_{2})+s_{1}-3\omega_{1}\\[1.99997pt] (s_{2}c_{1}-c_{2}s_{1})+s_{2}-3\omega_{2}\\[1.99997pt] s_{1}^{2}+c_{1}^{2}-1\\[1.99997pt] s_{2}^{2}+c_{2}^{2}-1\end{pmatrix}$ (4.4)
@@
Figure 4: Routing points and connecting paths for the discriminant of the Kuramoto model with three oscillators (4.4). The right picture shows a zoomed-in version of the left picture. The gradient flow is visible in blue, with arrows pointing towards index-zero routing points. We see that 77 of the $9$ regions are within the star-shaped figure. The last two regions are the inside of the ellipse and the unbounded region outside the ellipse. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
@@
To model its motion, two coordinate frames are used: a fixed frame attached to the ground, where the anchor points lie at fixed positions, and a moving frame attached to the triangular platform, where the platform’s corner points remain fixed relative to each other. In Figure 5, the leg lengths are $\ell_{1},\ell_{2}$, and $\ell_{3}$, and we work with their squares $c_{i}=\ell_{i}^{2}$ for convenience. In the fixed frame, the anchor points are located at $(0,0)$, $(A_{2},0)$, and $(A_{3},B_{3})$ while in the moving frame, the platform corners are located at the points $P_{1}=(0,0)$, $P_{2}=(a_{2},0)$, and $P_{3}=(a_{3},b_{3})$.
@@
A fundamental problem is to determine all the ways a real motion of the mechanism can start and end in the same “home” configuration for a fixed set of leg lengths. The configuration variables are $(p,\phi)=(p_{1},p_{2},\phi_{1},\phi_{2})$, where $p=(p_{1},p_{2})$ represents the translation of the platform and $\phi=(\phi_{1},\phi_{2})$ is a unit circle representation of its rotation. This yields a polynomial system $G(a,A,b_{3},B_{3},c;p,\phi)$ with variables $(p,\phi)$, parameters
@@
Figure 6: Routing points and connecting paths for the two-parameter version of the 3RPR system. The right picture shows a zoomed-in version of the left picture. The gradient flow is visible in blue. The arrows point towards index-zero routing points. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
@@
Finally, we consider another three-parameter version of the problem, where we instead view $(c_{1},c_{2},A_{2})$ as free parameters, and fix other parameters as in (4.5). In this case, we have not been able to find a symbolic expression for the discriminant surface, but its degree can readily be determined to 24 through a pseudo-witness set.
This example is at the limit of what our current implementation can handle, as the monodromy step did not terminate after 14 days [^1] Computations were run on a 4-socket Intel Xeon Gold 6128 system (24 cores total, 3.4 GHz) with 3 TB RAM.. At this point, 457 complex critical points for the routing function centered at $(4.72,4.33,1.70)$ had been found, of which 90 points were real. While likely not being a complete set of routing points, these points still provide a valuable sample of the parameter space. In particular, we can use them to verify that there are open regions in parameter space with the real root counts 2, 4 and 6, respectively.
@@ 4.3 Allee effect
To this end, we extend the discriminant $\Delta$ from (4.1) to $\Delta\cup V(ab(b-\frac{1}{2}))$. As explained in Remark 3.1, on the level of routing functions, this means adding $\log(|a|)$, $\log(|b|)$ and $\log(|b-\frac{1}{2}|)$ to $\log(r)$. It is clear from inspection of $G$ that $z=0$ is always a solution for biologically relevant parameters, and that all other nonnegative solutions have strictly positive coordinates. Hence, the number of nonnegative solutions is constant in each region of $((0,\infty)\times(0,\frac{1}{2}))\smallsetminus(\Delta\cup V(ab(b-\frac{1}{2})))$.
@@
An interesting property of this example is that $\Delta$ is not irreducible, and that some of its irreducible components have a non-reduced pseudo-witness set with respect to the equations in  (4.1). Hence, if these equations are used, our implementation effectively works with the subvariety $\hat{\Delta}\subset\Delta$ corresponding to the reduced part (see Remark 3.11). We can see the effect of this in the left part of Figure 7, which shows the resulting routing function, the biologically relevant routing points and the number of nonnegative solutions of $G(a,b,z)=0$ inside the green critical points. As we can see, gradient flow connects a parameter pair $(a,b)$, such that $G(a,b,z)=0$ has 15 positive solutions with another pair, for which there are 9 positive solutions.
Figure 7: Contours and critical points of a routing function for the modified discriminant $\Delta\cup V(ab(b-\frac{1}{2}))$ of the Allee system from Section 4.3. The left figure corresponds to using the formulation (4.1), which gives the subvariety $\hat{\Delta}$. The right figure uses the radical and gives the full discriminant $\Delta$. The routing points are in green and are labeled by the number of nonnegative solutions of $G(a,b,z)=0$. In both figures, there is a routing point of the unbounded region which is outside the picture, with nonnegative root count 3. Contours (up to a constant), critical points and gradient flow were computed without access to the discriminant polynomial.
@@
$I=\left\langle G(a,b,z),\ \det(J_{z}G(a,b,z))\right\rangle$
can readily be computed symbolically in the computer algebra system OSCAR [44, 18]. By using a generating set of $\sqrt{I}$, together with the standard “squaring up” technique [4, Section 2.5] for dealing with overdetermined systems, we obtain a reduced pseudo-witness set for $\Delta$. The resulting routing function and the biologically relevant routing points are shown in the right part of Figure 7, which agrees with the regions displayed in [45, Figure 3].